OpenStudy (anonymous):

Can someone tell me what the discriminant is in this quadratic equation. -x2 – 2x+5=0 -x2 – 2+5=0 x= (-b±√(b^2-4ac))/2a

OpenStudy (whpalmer4):

if the polynomial is in the form \(ax^2+bx+c=0\) the discriminant is \(b^2-4ac\) you have \(a=-1, b=-2, c=5\) discriminant is \((-2)^2-4(-1)(5)=\)

OpenStudy (anonymous):

Thank you Whpalmer4. I understood that the discriminant is the number of roots but I got lost in my equation solving it. Thanks you! Here is your medal!

OpenStudy (whpalmer4):

oh, it's not the number of roots! but the value will tell you something about the kinds of roots you will get...

OpenStudy (whpalmer4):

discriminant > 0 means 2 real roots discriminant = 0 means 2 identical real roots discriminant < 0 means 2 complex roots (this is true for a quadratic, I make no promises about the discriminant for higher order polynomials)

OpenStudy (anonymous):

ok, then what is the value of the roots in this equations?

OpenStudy (whpalmer4):

the value of the roots will be \[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a} = \frac{-(-2)\pm\sqrt{(-2)^2-4(-1)(5)}}{2(-1) }= \]

OpenStudy (whpalmer4):

\[=\frac{-2\pm\sqrt{24}} {-2}= -1\pm\sqrt{6}\]

OpenStudy (anonymous):

Ok, when I plugged in the discriminant I got 24 as the result; but the way that I solved my equation is differs a bit. hmmmm

OpenStudy (whpalmer4):

It's okay if the way you solved it differs a bit. What's not okay is if you got different answers :-)

OpenStudy (whpalmer4):

did you solve it by completing the square? \[-x^2-2x+5=0\]\[x^2+2x=5\]\[(x+1)^2 = 5+1^2\]\[x+1=\pm\sqrt{6}\]\[x=-1\pm\sqrt{6}\]

OpenStudy (anonymous):

Yes I got the square root of 8 over 2.

OpenStudy (whpalmer4):

for the equation \(-x^2-2x+5=0\)?

OpenStudy (anonymous):

Square root (8) =2.83

OpenStudy (whpalmer4):

Let's check that out: \[-(\frac{\sqrt{8}}{2})^2-2(\frac{\sqrt{8}}{2})+5=0\]\[-\frac{8}{4}-\sqrt{8}+5=0\]Uh....not going to end well...

OpenStudy (anonymous):

Ok, that means that I made an error somewhere

OpenStudy (whpalmer4):

Let's try mine: \[(-1-\sqrt{6})^2 = (1-\sqrt{6}-\sqrt{6}+(\sqrt{6})^2) = 7+2\sqrt{6}\] \[-(7+2\sqrt{6}) - 2(-1-\sqrt{6}) + 5 = 0\]\[-7-2\sqrt{6}+2+2\sqrt{6} + 5= 0\]\[-7 + 2 + 5 = 0\checkmark\]

OpenStudy (anonymous):

How do I check to see if -7+2+5=0 is equal to my equation -2-2+5=0

OpenStudy (whpalmer4):

Well, the point is that when I put my solution into the original equation, the polynomial evaluated to 0, just like it is supposed to: \(-x^2-2x+5 = 0\) \[-2-2+5=0\]\[-4+5=0\]\[1=0\]That's not true, is it? So if that's the result of putting your solution into the original equation, you've made a mistake either in finding the solution, or in testing it. Or maybe both :-)

OpenStudy (whpalmer4):

But any valid solution to the equation will give you a true statement if substituted into the original equation you solved.

OpenStudy (anonymous):

You are right, I tried to take a shortcut to find the discriminant and did not list the original equation: x^2 -3x + 5= 0

OpenStudy (whpalmer4):

\(x^2\) or \(-x^2\)? Without the - sign, you'll have complex numbers for your solution, are you expecting that?

OpenStudy (anonymous):

Yes, that is why I ended with the sqrt.(8)/2 man! Can you tell me what my discriminant is for this equation?

OpenStudy (whpalmer4):

is it -x^2 or +x^2? You say one thing here, and another in the PM

OpenStudy (anonymous):

Sorry about the mix up, I was writing from my notes and wanted to make sure I wrote my equation out correctly for class! I am straight now!

OpenStudy (whpalmer4):

but I'm not! which one is it?

OpenStudy (anonymous):

My final result is (8) = 2.83 or 0 + (-1.42) = -1.42 I find that the parabola crosses the x-axis also known as the x-intercepts.

OpenStudy (whpalmer4):

For WHICH EQUATION?!?

OpenStudy (anonymous):

the original equation which is -x^2 -3x+5=0

OpenStudy (anonymous):

After dinner I will watch some media on Quadratic formulas and hopefully that will help!

OpenStudy (whpalmer4):

Okay, how are you solving your quadratic equations? By the formula, factoring, or completing the square?

OpenStudy (whpalmer4):

Solutions to \[-x^2-3x+5=0\] do not have an 8 in them.

OpenStudy (whpalmer4):

If using the formula: \[a = -1, b = -3, c = 5\]\[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a} = \frac{-(-3)\pm\sqrt{(-3)^2-4(-1)(5)}}{2(-1)} \]\[= \frac{3\pm\sqrt{9+20}}{-2} = \frac{1}{2}(-3\pm\sqrt{29}) \]

OpenStudy (anonymous):

When we simplify o + sqrt: 0-4 (-1) (2) = 0+ sqrt 8/-2 = sqrt -8/2 = sqrt 8/2

OpenStudy (whpalmer4):

Where, oh where, are you finding a \(\sqrt{8}\)? Can you show me your work?

OpenStudy (anonymous):

Is there a way that I can e-mail my Word document so that you can see it in the quadratic formula?